OpenStudy (anonymous):

Which equation would best help solve the following problem? Rita hits a tennis ball with an initial velocity of 15 meters/sec. When she hits the ball, it is one meter above the ground. How long will it take for the ball to hit the ground? A. 4.9t2 + 15t + 1 = 0 B. –16t2+ 15t + 1 = 0 C. –4.9t2 + 15t = 1 D. –4.9t2 + 15t + 1 = 0

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@.Sam.

OpenStudy (sidsiddhartha):

use h = ut + gt^2 / 2 h = 1m, u = 15 and g = -9.8. so -4.9 t^2 + 15t = 1

OpenStudy (anonymous):

@Preetha

OpenStudy (sidsiddhartha):

C

OpenStudy (anonymous):

thank you